Solve problems with two unknown
A telephone cable 98 meters in length is cut into two pieces. If one piece is 22 meters longer than the other, how long are the two pieces?
Thank you.. I want to learn this subject, so I need a clear explanation

Answer by ReadingBoosters(3246)

The 98 m cable is cut into two pieces:
let
x = longer piece
y = shorter piece
so
x = y + 22 (it is 22 m longer)
x+y = 98 (the total length)
substitute x into the second equation
(y+22) + y = 98
Combine y
2y + 22 = 98
Subtract 22 from both sides to isolate y
2y + 22 - 22 = 98 - 22
2y = 76
Divide both sides by 2 to solve for y
y = 38 m is the length of the shorter piece
If x = y+22, then x=38+22 = 60m is the length of the longer piece
...
Proof
60 + 38 = 98
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.